### v2

# Realistic binary neutron star initial data with Elliptica

Abstract
This work introduces the Elliptica pseudo-spectral code for generating initial data of binary neutron star systems. Building upon the recent Elliptica code update, we can now construct initial data using not only piecewise polytropic equations of state, but also tabulated equations of state for these binary systems. Furthermore, the code allows us to endow neutron stars within the binary system with spins. These spins can have a magnitude close to the mass shedding limit and can point in any direction.
Keywords: initial data, binary neutron star, elliptic solver, equation of state
1 Introduction
Binary Neutron Star (BNS) systems are very common in our Universe. Estimates suggest their merger rate falls between $250$ to $2810$ Gpc-3yr-1 [1]. The coalescence of these BNS systems is a source of myriad phenomena such as, among others, ejecta [2, 3], accretion disk [4], jets [5], r-process nucleosynthesis [6], and kilonova [7]. These phenomena are treasure trove of information revealing aspects of physics at large scales, for instance, gravity in strong regimes and physical constants of the Universe, as well as aspects of physics regarding small scales, like, Equation of State (EOS) in supranuclear dense matter and formation of heavy elements in the periodic table.
As such, current detectors such as LIGO [8], VIRGO [9], and KAGRA [10] and the next generation ones like Cosmic Explorer [11], the DECi-hertz Interferometer Gravitational-wave Observatory (DECIGO) [12], Einstein Telescope [13], LIGO Voyager [14], the Laser Interferometer Space Antenna (LISA) [15], NEMO [16], and TianQin [17] are designed to look eagerly into sky and observe the physical signals emitted from the coalescence of compact binaries.
To unlock the wealth of information encoded in gravitational waves and their electromagnetic counterparts, accurate theoretical models are crucial. These models are essential for understanding events like GW170817 [18], the short gamma-ray burst GRB170817A [19], and the kilonova transient AT2017gfo [20]. However, finding analytical solution of the governing Partial Differential Equations, when BNS systems are coalescing, is not feasible as the PDE are in a highly non-linear regime where no approximation is applicable [21]. In light of this, Numerical Relativity (NR) community have put significant efforts to solve these equations numerically and hence make sense of the observations.
Simulation of compact binary system in NR often involves two steps. The first step is to find the solution of Einstein-Euler PDE on a hypersurface of the spacetime manifold, namely, constructing constraint satisfying and self-consistent Initial Data (ID) that present the binary system of interest at some time. ID codes such as, COCAL [22, 23, 24], Elliptica [25], FUKA [26], LORENE [27, 28, 29, 30, 31], NRPyElliptic [32], SGRID [33, 34, 35], SpECTRE’s elliptic solver [36, 37], Spells [38, 39, 40, 41], TwoPunctures [42, 43], are developed and utilized for this step.
The next step is to use the ID and simulate the system’s evolution over time; hence we can find the solution over the spacetime of interest. Dynamical evolution codes, for instance, AthenaK [44, 45], BAM [46, 47, 48], BAMPS [49, 50, 51], Dendro-GR [52], Einstein Toolkit [53, 54], ExaHyPE [55] GR-Athena++ [56, 57, 58] GRaM-X [59, 60], GRChombo [61, 62], Nmesh [63], NRPy+ [64], SACRA-MPI [65], Simflowny [66], SpEC [67, 68], SpECTRE [69, 70], and SPHINCS_BSSN [71], among others, are made and employed for this step.
Previously, Elliptica [25] was limited to the construction of ID for black hole neutron star binary systems. Additionally, the code was only supporting polytropic or piecewise polytropic EOS. In this work, we extend Elliptica’s infrastructure to construction ID for BNS systems as well as supporting tabulated EOS, for instance, the CompOSE tables [72, 73, 74].
To this aim, the Einstein-Euler equations are cast into coupled non-linear elliptic PDE [34, 75, 76], and solved iteratively using Newton-Raphson method by the Schur domain decomposition [25]. During the solve, the cubed spherical patches adapt to the surface of the Neutron Stars thus separating matter and vacuum and preventing Gibbs phenomena in the spectral method. To achieve the target values of interest such as momenta of the system, mass and center of NS, they are checked and adjusted during the solve. The EOS are approximated by either (piecewise) polytropic or tabulate ones.
The remainder of paper is organized as follows. In Sec. (2) we explain the mathematical background and formalism we use for ID of BNS systems, in particular, Einstein-Euler equations and EOS in Elliptica. Sec. (3) details the underling algorithms and numerical techniques for construction of BNS’s ID. We present the implementation of tabulated EOS by spline interpolant means and diagnostics in Elliptica. Additionally, we explain the iterative procedure for finding the physical and constraint satisfying ID for BNS systems. In Sec. (4), we present various convergence tests and comparison against post Newtonian answers to showcase the proof of concept for the new version the code. In Sec. (5), we discuss the possible improvements and future work.
Throughout the article, we use geometric units where $G=1$ is the constant of gravity, $c=1$ is the speed of light, and solar mass $M_{\odot}=1$. Additionally, following [77], we absorb the speed of light $c$ in the definition of pressure and total energy density.
2 Formalism
2.1 Einstein-Euler equations
To derive the equations that govern the gravity and matter on a spatial-like hypersurface $\Sigma_{t}$ of a 4-dimensional manifold $\mathcal{M}$ with a metric $g_{\mu\nu}$, we first write the line element of $\mathcal{M}$ as
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-\alpha^{2}dt^{2}+\gamma_{ij}(dx^{i}+\beta^{i}dt)(dx^{j}+\beta^{j}dt).$ (1)
This specific form of the line element is particularly well-suited for the $3+1$ formalism. It provides a clear view of the key variables (fields): $\alpha$, $\beta^{i}$, and $\gamma_{ij}$. Here, $\alpha$ is the lapse gauge and indicates the way that a sequence of spatial slices, $\Sigma_{t}$s, are combined to form the complete spacetime manifold $M$. $\gamma_{ij}$ is the induced 3-metric on each spatial hypersurface and can be written as $\gamma_{\mu\nu}=g_{\mu\nu}+n_{\mu}n_{\nu}$, where $n^{\mu}$ is the normal vector on $\Sigma_{t}$. $\beta^{i}$ is the shift vector. The importance of this gauge, i.e., shift vector, is that it represents the coordinate frame being used in each hypersurface. In this work, we take the shift vector as follows
$\beta^{i}=B^{i}+\epsilon_{ijk}{\Omega}^{j}_{\mathrm{BNS}}(r^{k}-r^{k}_{\mathrm{CM}})+\frac{v_{r}}{r_{\mathrm{BNS}}}(r^{i}-r^{i}_{\mathrm{CM}}).$ (2)
Here, $\Omega^{i}_{\mathrm{BNS}}$ is the orbital angular velocity of the BNS system, $r^{i}_{\mathrm{CM}}$ is the position of the system’s center of mass, $r_{\mathrm{BNS}}$ is the coordinate distance between the NS centers, $v_{r}$ is the radial velocity of the inspiraling coordinate system, and $\epsilon_{ijk}$ is the Levi-Civita symbol. Eq. (2) proves numerically convenient when applying Boundary Condition (BC) for $\beta^{i}$ at the edge of the computational grid where the position vector $\vec{r}$ has large values.
By definition the extrinsic curvature on $\Sigma_{t}$ is $K_{\mu\nu}=-\frac{1}{2}\pounds_{n}\gamma_{\mu\nu},$ in which £$\pounds_{n}$ is the Lie derivative along the normal vector. We note that by construction $K_{\mu\nu}n^{\mu}=0$, therefore, we can use spatial indices to describe the extrinsic curvature. Next, by utilizing conformal decomposition, we write
$\gamma_{ij}=$ $\psi^{4}\bar{\gamma}_{ij},$ (3)
$K^{ij}=$ $A^{ij}+\frac{1}{3}K\gamma^{ij}.$ (4)
Here $\psi$ is the conformal factor, $\bar{\gamma}_{ij}$ the conformal 3-metric, $A^{ij}$ the traceless part of $K^{ij}$. Moreover, since we are using the Extended Conformal Thin Sandwich (XCTS) formalism [75, 76]
$A^{ij}$ is decomposed as
$A^{ij}$ $=\psi^{-10}\bar{A}^{ij},$ (5)
$\bar{A}^{ij}$ $=\frac{1}{2\bar{\alpha}}\left((\bar{L}\beta)^{ij}-\bar{\gamma}^{ik}\bar{\gamma}^{jl}\bar{u}_{kl}\right),$ (6)
where
$\bar{u}_{ij}$ $=\frac{\partial\bar{\gamma}_{ij}}{\partial t},$ (7)
$(\bar{L}\beta)^{ij}$ $=\bar{D}^{i}\beta^{j}+\bar{D}^{j}\beta^{i}-\frac{2}{3}\bar{\gamma}^{ij}\bar{D}_{k}\beta^{k},$ (8)
$\alpha$ $=\psi^{6}\bar{\alpha},$ (9)
and $\bar{D}$ is the covariant derivative compatible with $\bar{\gamma}_{ij}$.
To complete the formulation we need to incorporate the source terms. We assume the fluid in NS are governed by ideal fluid, hence the stress energy tensor can be written as
$T_{\mu\nu}$ $=$ $(\rho_{0}+\rho_{0}\epsilon+P)u_{\mu}u_{\nu}+Pg_{\mu\nu},$
$=$ $\rho_{0}hu_{\mu}u_{\nu}+Pg_{\mu\nu},$
where, $\rho_{0}$ is the rest mass density, $\epsilon$ the specific internal energy, $P$ the pressure, $h$ the specific enthalpy, and $u^{\mu}$ the 4-velocity of the fluid. Additionally, for $3+1$ decomposition purposes, we project the stress energy tensor with respect to $\Sigma_{t}$ as follows
$E$ $=n_{\mu}n_{\nu}T^{\mu\nu},$ (11)
$S$ $=\gamma^{ij}\gamma_{i\mu}\gamma_{j\nu}T^{\mu\nu},$ (12)
$j^{i}$ $=-\gamma^{i}_{\mu}n_{\nu}T^{\mu\nu},$ (13)
where, $E$ is the measured energy by the Eulerian observer whose 4-velocity is $n^{\mu}$. $S$ is the trace of matter stress tensor, and $j^{i}$ is the momentum flux.
Unique answer to a linear elliptic equation with a source is guaranteed by the maximum/minimum principle. This principle becomes important during solve of constraint equations for high mass NS (for further discussion see [78]). To maintain the maximum/minimum principle, we rescale the stress energy projections as
$E$ $=\psi^{-6}\bar{E},$ (14)
$S$ $=\psi^{-6}\bar{S},$ (15)
$j^{i}$ $=\psi^{-6}\bar{j}^{i}.$ (16)
Finally, following XCTS formalism, we write Einstein’s equations in quasi equilibrium condition
$\bar{D}^{2}\psi-\frac{1}{8}\psi\bar{R}-\frac{1}{12}\psi^{5}K^{2}+\frac{1}{8}\psi^{-7}\bar{A}_{ij}\bar{A}^{ij}+2\pi\psi^{-1}\bar{E}$ $=$ $0,$ (17)
$\bar{D}^{2}(\bar{\alpha}\psi^{7})-(\bar{\alpha}\psi^{7})\left[\frac{1}{8}\bar{R}+\frac{5}{12}\psi^{4}K^{2}+\frac{7}{8}\psi^{-8}\bar{A}_{ij}\bar{A}^{ij}\right]$
$+\psi^{5}(\partial_{t}K-\beta^{k}\partial_{k}K)-2\pi\bar{\alpha}\psi^{5}(\bar{E}+2\bar{S})$ $=$ $0,$ (18)
$2\bar{\alpha}\left[\bar{D}_{j}(\frac{1}{2\bar{\alpha}}(\bar{L}\beta)^{ij})-\bar{D}_{j}(\frac{1}{2\bar{\alpha}}\bar{u}^{ij})-\frac{2}{3}\psi^{6}\bar{D}^{i}K\right]-16\pi\bar{\alpha}\psi^{4}\bar{j}^{i}$ $=$ $0,$ (19)
with the BC
$\lim_{r\to\infty}\psi=1,\ \ \ \lim_{r\to\infty}B^{i}_{0}=0,\ \ \ \lim_{r\to\infty}\alpha\psi=1.$ (20)
Additionally, we pick the free data as
$\bar{\gamma}_{ij}$ $=\delta_{ij},$ (21)
$K$ $=0,$ (22)
$\bar{u}_{ij}$ $=0.$ (23)
For hydrodynamic equations, following [34], we decompose the fluid into two parts: the rotational part of the fluid and is represented by a cross product, and the irrotational part of the fluid that is represented by a velocity potential (see, e.g.,[79]). In particular, the rotational part, which represents the NS spin, reads
$w^{i}=\epsilon_{ijk}\Omega^{j}_{\mathrm{NS}}(x^{k}-x^{k}_{c}),$ (24)
here, $\Omega^{j}_{\mathrm{NS}}$ is a free parameter to set the spin level, see Sec. (4.3), and $x^{k}_{c}$ denotes the NS coordinate. The irrotational part is shown by the potential $\phi(x,y,z)$ and obeys the following equations. [35]
$\frac{c\left(\rho_{0}\right)\alpha}{h}\psi^{-4}\bar{\gamma}^{ij}\partial_{i}\partial_{j}\phi-\frac{\rho_{0}\alpha}{h}\psi^{-4}\bar{\gamma}^{ij}\bar{\Gamma}^{k}_{ij}\partial_{k}\phi+2\frac{\rho_{0}\alpha}{h}\psi^{-5}\bar{\gamma}^{ij}(\partial_{i}\psi)(\partial_{j}\phi)$
$+\left(D_{i}\frac{\rho_{0}\alpha}{h}\right)\left(D^{i}\phi\right)D_{i}\left[\frac{\rho_{0}\alpha}{h}w^{i}-\rho_{0}\alpha u^{0}(\beta^{i}+\xi^{i})\right]=0,$ (25)
where, $\partial_{i}$ denotes the spatial partial derivative with respect to Cartesian coordinate $x^{i}$ and $\xi^{i}$ is the Killing vector. Here, following the smoothing method in [35]
$c(\rho_{0})=\rho_{0}+\bar{\epsilon}\rho_{0c}\left(\frac{\rho_{0c}-\rho_{0}}{\rho_{0c}}\right)^{4},$ (26)
in which, $\rho_{0c}$ is the values of $\rho_{0}$ at the NS center, and $\bar{\epsilon}$ is a constant number – for which we generally use $0.1$. Since the fluid is enclosed inside the patches that cover the NS, we further need to impose a BC on the NS surface as
$D^{i}\phi D_{i}\rho_{0}+w^{i}D_{i}\rho_{0}-hu^{0}(\beta^{i}+\xi^{i})D_{i}\rho_{0}=0.$ (27)
We note that since Eqs. (25,27) only include derivatives of $\phi$, we can not uniquely determine $\phi$. Hence, we further demand that the value of $\phi$ at the NS center is a constant number – like 0.
2.2 Equation of state
In order to close the system of equations we need an EOS. Elliptica deploys specific enthalpy, $h$, to create a link between the macroscopic properties of the matter and the gravity. As such, if we have a piecewise EOS we write
$\rho_{0}(h)$ $=K_{i}^{-n_{i}}\left(\frac{h-1-a_{i}}{n_{i}+1}\right)^{n_{i}},$
$P(h)$ $=K_{i}^{-n_{i}}\left(\frac{h-1-a_{i}}{n_{i}+1}\right)^{n_{i}+1},$
$\epsilon(h)$ $=\frac{a_{i}+n_{i}(h-1)}{n_{i}+1},$ (28)
where $n_{i}=\frac{1}{\Gamma_{i}-1}$ is the polytropic index and $K_{i}$’s are specific to the given EOS; $a_{i}$’s ensure the continuity of EOS [77] and are set as
$a_{0}$ $=0,$
$a_{i}$ $=a_{i-1}+\frac{K_{i-1}}{\Gamma_{i-1}-1}\rho_{i}^{\Gamma_{i-1}-1}-\frac{K_{i}}{\Gamma_{i}-1}\rho_{i}^{\Gamma_{i}-1}.$ (29)
Similarly, for tabulated EOS we represent $\rho_{0}\left(h\right)$, $P\left(h\right)$, and $\epsilon\left(h\right)$ as spline interpolants given discrete data points $\left(\rho_{0},P,\epsilon,h\right)$. Although typically only the quantities $\left(\rho_{0},P,\epsilon\right)$ or $\left(\rho_{0},P,e\right)$ are provided in tables, where $e$ is the total energy density, we may convert between variables via the relations:
$e$ $=\rho_{0}\left(1+\epsilon\right),$
$h$ $=1+\epsilon+\frac{P}{\rho_{0}}.$ (30)
Having written thermodynamic variables in terms of specific enthalpy, we now write specific enthalpy in terms of the metric and fluid variables [34]
$h$ $=$ $\sqrt{L^{2}-(D_{i}\phi+w_{i})(D^{i}\phi+w^{i})},$
$L^{2}$ $=$ $\frac{b+\sqrt{b^{2}-4\alpha^{4}[(D_{i}\phi+w_{i})w^{i}]^{2}}}{2\alpha^{2}},$
$b$ $=$ $[(\xi^{i}+\beta^{i})D_{i}\phi-C]^{2}+2\alpha^{2}(D_{i}\phi+w_{i})w^{i}.$ (31)
3 Numerical method
3.1 Coordinate system
The computation grid is tiled by cubed spherical coordinate systems [80], except at the NS centers, where we use simple Cartesian boxes to avoid coordinate singularities at $r=0$ [25, 81]. The relation between Cartesian coordinate, denoted by $x^{i}=(x,y,z)$, and cubed spherical coordinate, $X^{i}=(X,Y,Z)$, is
$X(x,y,z)=\frac{x}{z},\;Y(x,y,z)=\frac{y}{z},\;Z(x,y,z)=\frac{z-r_{\mathrm{in}}}{r_{\mathrm{out}}-r_{\mathrm{in}}},$ (32)
here, $X$ and $Y$ take value $\in[-1,1]$, and $Z\in[0,1]$. Additionally, $r_{\mathrm{in}}$ and $r_{\mathrm{out}}$ are defined
$r_{\mathrm{in}}$ $=\frac{\sigma_{\mathrm{in}}(X,Y)}{\sqrt{1+X^{2}+Y^{2}}},\;r_{\mathrm{out}}$ $=\frac{\sigma_{\mathrm{out}}(X,Y)}{\sqrt{1+X^{2}+Y^{2}}}.$ (33)
where, the shape of the inner boundary of a patch along the radial direction is determined by $\sigma_{\mathrm{in}}(X,Y)$ and its outer boundary by $\sigma_{\mathrm{out}}(X,Y)$. $\sigma(X,Y)$ is related to Cartesian coordinates by the equation $\sigma(X,Y)=\sqrt{x^{2}+y^{2}+z^{2}}$. Finally, we note that while Eq. (32) is written for patches along the z-axis, one can generalize this along any other axes.
As the distance from NS increases, we expect the fields fall as powers of $r^{-1}$. To account for this behavior we use a new transformation for $Z$ coordinate, denoted by $\widetilde{Z}$, for the paches covering large radii of the grid. The transformation reads
$\widetilde{Z}=\frac{\sigma_{\mathrm{out}}}{\sigma_{\mathrm{out}}-\sigma_{\mathrm{in}}}\left(1-\frac{\sigma_{\mathrm{in}}}{r}\right),$ (34)
Here $r=\sqrt{x^{2}+y^{2}+z^{2}}$ and still $\widetilde{Z}\in[0,1]$.
Finally, Elliptica uses Chebyshev polynomials of the first kind for the basis of the spectral expansion and deploys the extrema of Chebyshev polynomials for its collocation points [25].
Figure 1: Grid patches for a BNS system. Left: shown is the $x-y$ plane of the computational grid that is covered by different patches. To fully represent the grid, the other boundary is set to a shorter length. For a production run, the outer boundary is set generally at a radius of $10^{5}\mathrm{\,M_{\odot}}$ from to the center of grid. Right: a zoomed-in view of the $x-y$ plane focusing on the NS regions. A Cartesian box is used around each NS center to avoid coordinate singularities. By using various $\sigma(X,Y)$ values, cubed spherical patches can adapt to different shapes, effectively capturing the shape of the NS surface. This allows for the treatment of matter and vacuum in separate patches.
Fig. (1) illustrates the patches covering the $x-y$ plane of the computational grid used in this work. It is worth noting that the sizes used for the grid and NS are reduced to illustrate the complete computational grid. In practice, we set the outer boundary at a radius of $10^{5}\mathrm{\,M_{\odot}}$.
3.2 Elliptic solver
Given an elliptic PDE, we linearize the equation to use the Newton-Raphson method (see [25] for a complete description). Thus, the original elliptic PDE becomes a matrix equation, like $Ax=B$, to be solved. Now the challenge is to solve this matrix equation efficiently. To this aim, we use the Schur domain decomposition method [25, 82]. At the core of this method there is a divide and conquer strategy; in this strategy, first the coupled equations are solved, and then the system decomposes into independent (uncoupled) subsystems, amenable for parallel solving. The Schur domain decomposition allows solving the full system using shared memory multiprocessing in which each CPU core is assigned to solve a matrix equations at each patch independently.
The Schur domain decomposition method arranges the Jacobian matrix in Newton-Raphson method into two parts. The first part is associated with all uncoupled equations and the second part comprises the coupled equations. These couplings between equations occur due to the inter-BC at the interfaces of adjacent patches [25]. As such, the system of matrix equations can be seen as two equations with two unknowns as follows
$\left(\begin{array}[]{c c}B&E\\ F&C\end{array}\right)\left(\begin{array}[]{c}\tilde{v}\\ \tilde{w}\\ \end{array}\right)=\left(\begin{array}[]{c}f\\ g\\ \end{array}\right),$ (35)
where, the vector $\tilde{v}$ denotes all the unknowns that are uncoupled, i.e., $\tilde{v}$ comprises all unknowns that stem from inner points of a patch and hence uncoupled from the other patches. The vector $\tilde{w}$ is the unknowns that are due to coupled equations, i.e., stem from inter-BC between adjacent patches. By such an arrangement, if the unknown $\tilde{w}$ is solved, then the equation of $\tilde{v}$ can be solved consequently.
Therefore, we first solve for $\tilde{w}$, and then we use it to solve for $\tilde{v}$. Accordingly, in Elliptica, we solve Eq. (35) like:
$(C-FE^{\prime})\tilde{w}$ $=g-Ff^{\prime},$ (36)
$\tilde{v}$ $=f^{\prime}-E^{\prime}\tilde{w},$ (37)
where
$E^{\prime}$ $=B^{-1}E,$
$f^{\prime}$ $=B^{-1}f.$ (38)
As we can see eq (36) only involves the unknown $\tilde{w}$, and it can be summarized as
$S\tilde{w}=g^{\prime},$ (39)
where, $S$ is called Schur complement matrix.
After solving Eq. (39) for $\tilde{w}$, we solve for $\tilde{v}$ using Eq. (37) – hence we find the solution of the whole system.
Algorithm 1 Schur complement domain decomposition method.
1:Solve $BE^{\prime}=E$ for $E^{\prime}$;
2:Solve $Bf^{\prime}=f$ for $f^{\prime}$;
3:Compute $g^{\prime}=g-Ff^{\prime}$;
4:Compute $S=(C-FE^{\prime})$;
5:Solve $S\tilde{w}=g^{\prime}$ for $\tilde{w}$;
6:Compute $\tilde{v}=f^{\prime}-E^{\prime}\tilde{w}$;
A summary of the Schur complement domain method to solve elliptic equations is shown in algorithm 1. For a more in-depth discussion about implementation of Schur domain decomposition and parallelization of matrix solver the reader may consult [25]. Finally, we use the publicly available and open-source UMFPACK direct solver [83].
3.3 Diagnostics
In Elliptica we compute the baryon mass of each NS using [84]
$M_{B}=\int_{\mathrm{NS}}\rho_{0}\alpha\psi^{6}\sqrt{\bar{\gamma}}d^{3}x,$ (40)
in which $\bar{\gamma}$ is the determinant of $\bar{\gamma}_{ij}$ and the integration is taken over the volume of the NS.
To measure the NS spins, we can use the flat space coordinate rotational Killing vector, following [85], on the surface of NS:
$\vec{\phi}_{x}$ $=-(z-z_{c})\vec{\partial}_{y}+(y-y_{c})\vec{\partial}_{z},$
$\vec{\phi}_{y}$ $=+(z-z_{c})\vec{\partial}_{x}+(x-x_{c})\vec{\partial}_{z},$
$\vec{\phi}_{z}$ $=-(y-y_{c})\vec{\partial}_{x}+(x-x_{c})\vec{\partial}_{y},$ (41)
in which $(x_{c},y_{c},z_{c})$ is the coordinate center of the NS and $(\vec{\partial}_{i})_{i\in\{x,y,z\}}$ are the basis vectors associated with the Cartesian coordinates. Accordingly, NS spin $S_{i}$, for the $i$ direction, is measured on the surface of NS by
[…]
where, $s^{k}$ is the normal vector on the NS surface. The dimensionless spin is defined
$\chi_{i}:=\frac{S_{i}}{M^{2}_{\mathrm{NS}}}.$ (43)
Additionally, following [81], we can first compute the angular momentum $J_{i}$, the center $R^{i}_{c}$, and linear momentum $P_{i}$ of the NS, and then calculate $S_{i}$ as follows
$S_{i}=J_{i}-\epsilon_{ijk}R^{j}_{c}P_{k}.$ (44)
Since the chosen free data $\bar{\gamma}_{ij}$ in Eq. (21) satisfies the quasi-isotropic gauge condition [84] and $K$ in Eq. (22) meets the asymptotic maximal gauge condition [84], the ADM linear momentums and angular momentums of the system are defined [84]
[…]
here,
$\vec{\phi}_{x}$ $=-(z-z_{\mathrm{CM}})\vec{\partial}_{y}+(y-y_{\mathrm{CM}})\vec{\partial}_{z},$
$\vec{\phi}_{y}$ $=+(z-z_{\mathrm{CM}})\vec{\partial}_{x}+(x-x_{\mathrm{CM}})\vec{\partial}_{z},$
$\vec{\phi}_{z}$ $=-(y-y_{\mathrm{CM}})\vec{\partial}_{x}+(x-x_{\mathrm{CM}})\vec{\partial}_{y}.$ (47)
Lastly, to calculate the total ADM mass of the system, we use [84]
$M_{\mathrm{ADM}}=\int_{\Sigma_{t}}\left[\psi^{5}E+\frac{1}{16\pi}\left(\bar{A}_{ij}\bar{A}^{ij}\psi^{-7}-{\bar{R}}\psi-\frac{2}{3}K^{2}\psi^{5}\right)\right]\sqrt{\bar{\gamma}}d^{3}x.$ (48)
3.4 Tabulated equations of state
We can import tabulated EOS from the CompOSE repository [72, 73, 74]. A general CompOSE table is first restricted to $T=0$. The baryon density $n_{b}$ is converted to the rest mass density $\rho_{0}$ via the neutron mass $m_{n}$: $\rho_{0}=m_{n}n_{b}$.
We numerically differentiate and interpolate the EOS table to generate a Hermite spline representation of the functions $\rho_{0}\left(h\right)$, $P\left(h\right)$, and $\epsilon\left(h\right)$ (see Sec. (A)). In practice, the logarithms of these quantities are actually used to generate the interpolants.
A number of slight modifications may be made to the EOS in preprocessing in order to make it more amenable to interpolation. These modifications are necessary for both the physical consistency of the EOS and to facilitate the convergence of the solution.
The first modification involves adjusting specific enthalpy at the NS surface for the EOS tables in which the specific enthalpy falls below $1$. To this aim, at the minimum pressure, which we consider occurring at the beginning of the table, we scale all $\rho_{0}$ points by the multiplicative constant $\eta=h_{\mathrm{min}}$ corresponding to the lowest value of $h$ in the table. Accordingly, $h$ points are scaled by the value $\eta^{-1}$, i.e:
$\rho_{0}\to\eta\rho_{0}$
$h=\frac{e+P}{\rho_{0}}\to\frac{e+P}{\eta\rho_{0}}=\eta^{-1}h.$ (49)
This alteration can be carried through to dynamical evolution codes by e.g. using the same EOS table or by scaling the rest-mass density, at the minimum pressure, by the same factor used in the preprocessing step.
In addition, several other features of tabular EOS may pose problems for both interpolation and the use of specific enthalpy as the independent thermodynamic variable. Among these problems is the presence of a region near the surface of the NS where both $\frac{de}{dh}$ and $\frac{d\rho_{0}}{dh}$ diverge. This region is common to many EOS with a ‘crust’, including both tabular EOS such as SFHo [86] and piecewise polytropics [87]. While this region is unavoidable if we wish to accurately represent the equation of state via specific enthalpy, cf. [88, 89], it also limits the accuracy of the solution within the NS (see Fig. (3)).
Another consideration is the spacing of the data points in the CompOSE table, which may be highly irregular especially for tables generated as a piecewise combination of different models. To re-grid the EOS, we numerically differentiate the table points using Fornberg’s method, explained in Sec. (B), and generate a low-degree Hermite interpolant, Sec. (A), which is sampled to produce new data points on an evenly-spaced grid. This method decreases oscillations when the new data points are themselves interpolated. Additionally, the interpolation of the EOS truncates the jumps in derivatives of the thermodynamic variables, i.e., $\frac{de}{dh}$ and $\frac{d\rho_{0}}{dh}$, that as we mentioned before, in some tables are not well-defined. This low-degree Hermite interpolant corresponds to approximating the derivatives as finite at these points (since infinite values would be unphysical anyway).
Figure 2: Comparison of changes with respect to the specific enthalpy ($h$) in the rest mass density ($\rho_{0}$) near the NS surface for three EOS: K96 (single polytrope), SLy (piecewise polytrope), and SFHo (table). A lower specific enthalpy corresponds to a point closer to the NS surface, with $h=1$ defining the surface itself. $\rho_{0}$ is in geometric unit and $h$ is dimensionless.
Fig. (2) shows the resultant tabulated SFHo EOS after re-griding using a low-degree Hermite interpolant with comparison to EOS of K96 and SLy. The K96 EOS is a single polytrope, corresponding to $K=96.7$ and $\Gamma=2$ [90]. SLy is an approximate piecewise polytrope of tabulated SLy4, and therefore quantities such as $\frac{d\rho_{0}}{dh}$ are not continuous in certain regions corresponding to transitions between piecewise segments [91]. SFHo, here restricted to the zero-temperature regime, shows similar discontinuities [86].
Examining the functions relevant to the ID generation, e.g. $\rho_{0}$ and $e$, we see that the single-polytrope model is considerably simpler and more amenable to numerical solution than the models with discontinuities. The difference is typically relevant to the surface of the NS (where $h=1$), as shown in the nearly vertical profile of $\frac{d\rho_{0}}{dh}$ in Fig. (2). The sharp changes in $\frac{d\rho_{0}}{dh}$ with the piecewise polytropic and tabular EOS are not perfectly captured by the grid, and thus they limit the convergence of the constraint violations (see Sec. (4.1)).
There are several potential ways to improve the convergence properties of the code in this area. For example, one could apply filtering to the spectral coefficients [92] of $\rho_{0}$ or implement a routine to identify the location of discontinuities, thereby creating patches where the discontinuity occurs at the boundary between those patches. Additionally, using spectral EOS [88, 89, 93], and imposing continuity at the piecewise polytrope EOS [87] have been shown to enhance the continuity of the solutions. The implementation and comparison of these methods are left for future work.
3.5 Initial data construction
Construction of ID often necessitates an iterative approach. This process involves progressively refining the ID until a suitable solution is achieved. However, rapid or uncontrolled updates of the fields from one step to the next, or lack of adjustments of NS masses or their centers can lead to a divergent answer and code crashes. Another challenge is to find the NS surface after each update, as the true surface is unknown a-priori, and then creating a new computational grid with patch’s surfaces adapted to the new NS surface.
We deploy the following iterative procedure to construct ID of a BNS system. In particular, we start from a coarse resolution grid and progressively refine the answer, while controlling diagnostics, until the Hamiltonian and momentum constraints, Eqs. (55) and (56), reach a plateau. Then, we increase the resolution and repeat this iteration until we achieve the level of desired accuracy.
Step 0: We superimpose two Tolman-Oppenheimer-Volko (TOV) star solutions as the initial guess of the fields $\{\psi,\alpha\psi,B^{i}\}$. To initialize the $\phi$ fields for each NS we use the approximation $\phi=-\Omega^{z}_{\mathrm{BNS}}(y_{\mathrm{NS}}-y_{\mathrm{CM}})x$.
Step 1: We solve the elliptic equations iteratively in this specific order: first Eq. (25) for the matter field, and then Eqs. (17), (18), and (19) for the metric fields. This approach has been found to enhance the solution’s stability, particularly when solving the matter field first. The order of solving the metric fields themselves, however, appears to be less critical. Moreover, during each iteration step, we focus on solving a single elliptic equation while keeping the other fields fixed. In essence, the fixed fields act as source terms influencing the equation being solved. Finally, within the Newton-Raphson iterative method, we perform only one update step per field, and then incorporating the newly solved value into the source terms for the next equation. In the following $\Xi$ denotes a field from the set $\{\phi,\psi,\alpha\psi,B^{i}\}$.
Step 2: As mentioned earlier, iterative solvers are often sensitive to prompt changes; code crashes can happen if a sudden update take place in the system. As such, we update the field solutions that are obtained from Step 1 in a relaxed fashion. To this aim, we use $\Xi=\lambda\Xi_{\mathrm{new}}+(1-\lambda)\Xi_{\mathrm{old}}$, in which $\Xi_{\mathrm{old}}$ denotes the solution before entering Step 1 and $\Xi_{\mathrm{new}}$ is the solution after exiting that step. $\lambda$ denotes the weight of update. We generally use $\lambda=0.2$. This choice of $\lambda$ value is proven to work for all experiments we have done.
Step 3: After updating all $\Xi$ fields, we see the baryon mass of each NS deviates, often by a few percents, from the target value. Additionally, for spinning NS, since at Step 0 we used TOV solution, and then we added the spin vector to the NS, it is not surprising to observe the baryon mass is different from the prescribed target value. Moreover, the starting resolution is often coarse, and again the baryon mass may change as we go to a higher resolution. If we do not account for these changes, the baryon mass deviates even further at later iterations and may lead to a code crash. To adjust the baryon mass, we note that $\rho_{0}=\rho_{0}(h)$, and as shown in Eq. (31) the specific enthalpy depends on a constant $C$, i.e., $h=h(C)$, which implies $\rho_{0}=\rho_{0}(C)$. Therefore by using a root finder, we find the value of $C$ in Eq. (40) such that the baryon mass remains the same as the target value.
Step 4: Since at Step 0 we began with a rough approximation of the solution, the linear ADM momentums are initially not zero. Furthermore, for asymmetric masses or spinning systems, we do not know in advance where is the exact position of the system center of mass – as one needs to consider the full general relativity effect to find it. Therefore, we iteratively find the system center of mass $\vec{r}_{\mathrm{CM}}$, by demanding the ADM momentum in each direction to be zero. The ADM momentum in $z$-direction proves consistently to be small, generally, $\frac{|P^{\infty}_{z}|}{M_{\mathrm{ADM}}}<10^{-11}$, hence we keep $z_{\mathrm{CM}}=0$, as the original value. However, often the initial value of ADM momenta in $x$ or $y$ direction is $\approx 10^{-5}M_{\mathrm{ADM}}$. Therefore, we adjust $x_{\mathrm{CM}}$ and $y_{\mathrm{CM}}$ as follows
$x_{\mathrm{CM,new}}$ $=x_{\mathrm{CM,old}}+\lambda\frac{P^{\infty}_{y}}{\Omega^{z}_{\mathrm{BNS}}M_{\mathrm{ADM}}},$
$y_{\mathrm{CM,new}}$ $=y_{\mathrm{CM,old}}-\lambda\frac{P^{\infty}_{x}}{\Omega^{z}_{\mathrm{BNS}}M_{\mathrm{ADM}}}.$ (50)
Here, $\lambda$ is generally chosen $0.2$. It is worth noting that $K_{ij}$ is a function of $\beta^{i}$, while $\beta^{i}$ through Eq. (2) is a function of $r^{i}_{\mathrm{CM}}$. Therefore, by adjusting the $r^{i}_{\mathrm{CM}}$, we influence $K_{ij}$ – and hence all other fields as they are coupled to another – and we can derive $P^{\infty}_{i}$ to zero iteratively.
Step 5: We update the stress energy tensor, in particular, we update the specific enthalpy for each NS in a relaxed fashion as $h=\lambda h_{\mathrm{new}}+(1-\lambda)h_{\mathrm{old}}$; $\lambda$ is usually set to $0.5$. Then, we use the new value of specific enthalpy to update $\rho_{0}(h)$, $\epsilon(h)$, and $P(h)$.
Step 6: Finding the orbital angular velocity, $\Omega^{z}_{\mathrm{BNS}}$, for a quasi-circular orbit demands full solution of general relativity. Since we start with a Newtonian approximation for this value we need to refine it. Following [94], so called the force balance method, we first compute $\partial_{i}\ln\Gamma$ at each NS center, where $\Gamma$ computed as
$\Gamma=\frac{\alpha u^{0}[1-(\beta^{i}+\xi^{i}+\frac{w^{i}}{hu^{0}})\frac{D_{i}\phi}{\alpha^{2}hu^{0}}-\frac{w_{i}w^{i}}{(\alpha^{2}hu^{0})^{2}}]}{\sqrt{1-(\beta^{i}+\xi^{i}+\frac{w^{i}}{hu^{0}})(\beta_{i}+\xi_{i}+\frac{w_{i}}{hu^{0}})\frac{1}{\alpha^{2}}}}.$ (51)
Then, we use a root finder to find $\Omega^{z}_{\mathrm{BNS}}$ such that
$\partial_{i}\ln[\alpha^{2}-(\beta^{i}+\xi^{i}+\frac{w^{i}}{hu^{0}})(\beta_{i}+\xi_{i}+\frac{w_{i}}{hu^{0}})]+2\partial_{i}\ln\Gamma=0,$ (52)
where, $\partial_{i}=\frac{\partial}{\partial x^{i}}$. Since in our setup the NS centers sit on y-axis, we compute Eq. (52) along this axis. We find $\Omega^{z}_{\mathrm{BNS}}$ for each NS centers and update it accordingly.
Step 7: At this point the specific enthalpy profile and hence NS surface are changed. We need to find the new location of NS surface for adjusting cubed spherical patches that are covering the NS – so we can separate matter and vacuum into different patches. Since some parts of the NS may need to extend to the patches that are currently covering vacuum, we extrapolate $h$ into these patches, so the root finder can find where $h=1$. To this aim we use the following formula to extrapolate specific enthalpy to the vacuum
$f\left(r\right)=\left(a+\frac{b}{r}\right)\exp\left(-c_{0}\frac{r}{r_{0}}\right),$ (53)
here, $r$ denotes the coordinate distance from the NS center, and $c_{0}$ is a constant (typically $0.01$). We find the values of $a$, and $b$ by demanding $C^{1}$ continuity across the NS surface. Additionally, we use the same Eq. (53) to extrapolate $\phi$ field outside the NS. This step is required when the NS surface is expanded, and we want to interpolate $\phi$ from the current grid to a new grid.
Step 8: We identify the center of each NS by locating the coordinate where the specific enthalpy reaches its maximum value. After multiple updates to the matter fields in previous steps, the NS centers can exhibit slight drifts from their initial positions. These drifts can accumulate over time, potentially leading to code crashes. To address this issue, we employ a corrective measure that adjusts the specific enthalpy function as follows
$h_{\mathrm{new}}(\vec{r})=h_{\mathrm{old}}(\vec{r})-(\vec{r}-\vec{r}_{0})\cdot\vec{\nabla}h_{\mathrm{old}}(\vec{r_{0}}),$ (54)
where, $r_{0}$ denotes the coordinate of NS center. This adjustment ensures that the maximum value of the specific enthalpy remains at the same location, effectively preventing the NS centers from drifting significantly.
Step 9: We find the profile of $\sigma_{\mathrm{out}}(X,Y)$ that is necessary for Eq. (33) to have patch surfaces fitting the NS surface. To this end, for a given angular $\theta$ and $\phi$ in spherical coordinate, we use a root finder for $r$ to solve $h(r,\theta,\phi)-1=0$ and hence finding the new NS surface. Here, $r$ is the coordinate distance from the NS center.
Step 10: If the NS surfaces are changed or if the resolution increases at the next iteration, we create a new grid. To ensure a smooth transition, we use spectral interpolation techniques to transfer data from the previous grid onto the new one.
Step 11: We monitor the Hamiltonian and momentum constraints, Eqs. (55) and (56). We restart from Step 1 unless the constraints reach their truncation error and are level off. In this case, when the constraints reach a plateau, we stop the iterative process if there is no higher resolution demanded; otherwise we go to the next resolution and start from Step 1.
4 Results
4.1 Convergence test
Since Elliptica is a pseudo-spectral code, the first expectation of the code is spectral convergence feature. As such, we calculate the Hamiltonian and momentum constraints, respectively, using
$H$ $:=R-K_{ij}K^{ij}+K^{2}-16\pi E$ (55)
$M^{i}$ $:=D_{j}(K^{ij}-\gamma^{ij}K)-8\pi j^{i}$ (56)
For the convergence test, we generate ID for two symmetric BNS systems: one with polytropic K96 EOS and the other with tabulated SFHo EOS. In these systems, the NS have baryon mass $1.4$ with no spin, and their separation is $50M_{\odot}$.
(a) K96, NS interior regions
(b) SFHo, NS interior regions
(c) K96, NS exterior regions
(d) SFHo, NS exterior regions
(e) K96, all regions
(f) SFHo, all regions
Figure 3: The $L_{2}$-norm of the Hamiltonian and momentum constraint violations for the BNS system with the K96 EOS (single polytrope) and the SFHo EOS (table). The constraint violations at each point are summed over all points in the specified regions. The regions not including a tabular EOS show exponential convergence, while regions with such an EOS reach a limit at high resolutions.
Fig. (3) shows the $L_{2}$-norm of the Hamiltonian and momentum constraints as a function of grid resolution (focusing on the final iteration at each resolution) for the two systems. In figs. (3(a),3(b)), the convergence test emphasizes on the NS. We observe the constraint violations decrease exponentially for smooth matter field, i.e., K96 EOS. For the SFHo EOS, initially for low resolutions the constraints converge exponentially but for higher resolutions the constraints level off. This behavior is expected; when the grid resolution is coarse the true discontinuities of EOS variables are not seen by the spectral expansion. For fine resolutions all features of EOS variables emerge and while spectral convergence try to resolve these features, but it is not successful. Hence, we see constraints are soon level off and do not decrease as we increase the resolution. We note that, for the tabular EOS (SFHo), the convergence is limited due to the sharp features of $\frac{d\rho_{0}}{dh}$ near the surface of the NS (see Fig. (2)). Indeed, the overall magnitude of the constraint violations is higher and the rate of convergence is not exponential with the tabular EOS (in contrast to the simple polytrope). Nonetheless, the profile of the convergence tests still suggest convergence up to a limit at high resolutions.
Ensuring spectral convergence of the code for regions that fields are smooth, we calculate $L_{2}$-norm of constraints at immediate neighboring regions of NS, where there is no matter fields. Figs. (3(c), 3(d)) demonstrates they convergence spectrally.
Finally, the overall converge of the constraints, calculated over all regions, are shown in figs. (3(e), 3(f)); while both systems exhibit convergence as the resolution increases, the rate of convergence for BNS with SFHo EOS is smaller than K96 EOS.
4.2 Post-Newtonian test
We calculate the binding energy $E_{b}$ of a symmetric non-spinning BNS system with NS baryon mass $1.4$ and tabulated SLy4 EOS for varied separations of NS. Here $E_{b}=M_{\mathrm{ADM}}-M_{\infty}$ and $M_{\infty}=M^{1}_{\mathrm{TOV}}+M^{2}_{\mathrm{TOV}}$; we use Eq. (48) to calculate $M_{\mathrm{ADM}}$ and $M^{1/2}_{\mathrm{TOV}}$ are the corresponding gravitational mass of NS in isolation – found by a TOV solver.
To determine $\Omega^{z}_{\mathrm{BNS}}$ for a specific separation, we employ an iterative approach based on the force balance method described in Eq. (52). To validate the accuracy of the generated ID against the expected analytical values for significant separations, particularly when the BNS system exhibits quasi-circular motion, we calculate the binding energy $E_{b}$ corresponding to the given $\Omega^{z}_{\mathrm{BNS}}$ utilizing a post-Newtonian formula, the 4 post-Newtonian order approximation, outlined in [95], Sec. 7.4. Specifically, to compute the post-Newtonian values of $E_{b}$, we need the reduced mass $\mu=M^{1}_{\mathrm{TOV}}M^{2}_{\mathrm{TOV}}/{M_{\infty}}$, the symmetric mass $\nu=\mu/M_{\infty}$, and the dimensionless orbital velocity $x^{3/2}=\Omega^{z}_{\mathrm{BNS}}M_{\infty}$.
Figure 4: Analytic post-Newtonian (PN) curve versus NR curve. The values of $E_{b}$ for given $\Omega^{z}_{\mathrm{BNS}}$ are compared for post-Newtonian method and NR method. For binaries with a large separation post-Newtonian and NR values are matching. For BNS of a close separation there is a slight deviation from post-Newtonian prediction as the system is not fully in quasi-circular status. Here, $\mu={M^{1}_{\mathrm{TOV}}M^{2}_{\mathrm{TOV}}}/{M_{\infty}}$ and the binary has a symmetric baryon mass of $1.4$ for each NS and uses tabulated SLy4 EOS.
Fig. (4) illustrates the binding energy of the BNS system across different separations compared to the post-Newtonian data points. In cases of binaries with significant separations, meeting the quasi-circular motion criteria, we observe that the post-Newtonian and NR values align. However, for smaller separations, the NR values of $E_{b}$ are marginally higher.
4.3 Spin
We generate ID for different $\Omega^{z}_{\mathrm{NS}}$, in Eq. (24), values pertaining to one of the NS in the BNS system, where each NS has a baryon mass of $1.4$, and they are separated by $30M_{\odot}$. In this scenario, we utilize the K96 EOS. The maximum dimensionless spin achievable is $\approx 0.56$, which corresponds to the mass shedding limit of a single NS as discussed in [96]. By increasing the $\Omega^{z}_{\mathrm{NS}}$ beyond $0.02$, the spinning NS becomes too oblate that the NS surface finding routine fails (Step 9 in Sec. (3.5)).
Figure 5: The relation between $\Omega^{z}_{\mathrm{NS}}$ and $\chi_{z}$ is depicted by the solid line for a symmetric mass BNS system with a baryonic mass of $1.4$ and a separation of $30\mathrm{\,M_{\odot}}$. The dashed line shows the corresponding values of $\chi_{z}$ for an isolated spinning NS for the same values of baryonic mass and EOS. We find an excellent agreement between spin values of the isolated NS and the BNS system. Initially, $\chi_{z}$ exhibits a linear growth pattern with $\Omega^{z}_{\mathrm{NS}}$, but as $\Omega^{z}_{\mathrm{NS}}$ is raised beyond a certain point, $\chi_{z}$ starts to increase non-linearly.
Additionally, to compare the spin values in a BNS system with those of an isolated spinning NS, we create ID for isolated NS with the same values of baryon mass, EOS, and $\Omega^{z}_{\mathrm{NS}}$ values (except the largest value  [^2] To create ID for $\Omega^{z}_{\mathrm{NS}}=0.02$, the NS surface finding routine fails, hence we can not report the corresponding spin value.). The resulting spin values are shown in Fig. (5), where we find excellent agreement between spin value of the isolated NS and the BNS system.
Fig. (5) shows the relation between $\Omega^{z}_{\mathrm{NS}}$ in Eq. (24) and $\chi_{z}$ in Eq. (43). There is a linear relation between $\Omega^{z}_{\mathrm{NS}}$ and $\chi_{z}$ for low values of $\Omega^{z}_{\mathrm{NS}}$. However, as $\Omega^{z}_{\mathrm{NS}}$ increases, the relationship transitions into a nonlinear pattern, causing $\chi_{z}$ to rise more steeply.
Moreover, Fig. (5) can be utilized as an approximate reference to determine the appropriate value of $\Omega^{z}_{\mathrm{NS}}$ corresponding to a desired spin level.
4.4 Outer boundary radius
We aim to study the effect of the outer boundary radius of the grid, where Eq. (20) is imposed, on the convergence behavior of the constraints and the binding energy of the system. To this end, we construct four ID sets for a symmetric non-spinning BNS system with NS baryon mass $1.4$ and the tabulated SLy4 EOS each featuring a different outer boundary radius. In particular, the chosen outer boundary radii are $5\times 10^{4}\mathrm{\,M_{\odot}}$, $1\times 10^{5}\mathrm{\,M_{\odot}}$, $5\times 10^{5}\mathrm{\,M_{\odot}}$, and $1\times 10^{6}\mathrm{\,M_{\odot}}$.
We find that setting the outer boundary radius larger than $10^{6}\mathrm{\,M_{\odot}}$, can lead to code failures due to issues with the patch finder routine in the code. This routine is essential for establishing the inter-BC [25] between adjacent patches, as it identifies neighboring patches for each patch. Consequently, additional code support would be necessary if a larger outer boundary is required. However, a grid with an outer boundary at $10^{6}\mathrm{\,M_{\odot}}$ is already few orders of magnitude larger than the typical grids used in evolution simulations, such as those in [57, 97, 98]. Therefore, this size is sufficiently large for the ID to be interpolated onto the grid of an evolution code.
Figure 6: Shown is the $L_{2}$-norm of Hamiltonian ($H$) and momentum ($M^{i}$) constraint violations for a symmetric BNS system with the tabulated SLy4 EOS and the outer boundary radius of $5\times 10^{5}\mathrm{\,M_{\odot}}$. The constraints are calculated for an arbitrary patch that covers large-radius regions of the grid, where the flat spacetime BC, Eq. (20), are imposed on its outermost boundary. Other patches within these large-radius regions display a similar convergence pattern.
Fig. (6) illustrates the convergence test of the Hamiltonian and momentum constraints focusing on one of the patches that cover the large-radius regions of the grid, i.e, reaching the outer boundary of the grid. We observe a clear spectral convergence for all constraints. This convergence behavior remains quantitatively unchanged across all other patches in these large-radius regions as well as across all other ID sets with different outer boundary radii. Furthermore, the convergence behavior for all regions, NS exterior regions, and NS interior regions are consistent with the findings for the SFHo EOS, and discussed in Sec. (4.1).
Figure 7: Relative difference in binding energy as a function of outer boundary radius. We define $\Delta E_{b}=\bar{E}_{b}-E_{b}$, in which $\bar{E}_{b}$ represents the binding energy of the BNS system with the largest outer boundary radius, i.e., $10^{6}\mathrm{\,M_{\odot}}$.
Additionally, in Fig. (7), we plot the absolute value of the relative difference in binding energy as a function of the outer boundary radius. We define $\Delta E_{b}=\bar{E}_{b}-E_{b}$, in which $\bar{E}_{b}$ represents the binding energy of the BNS system with the largest outer boundary radius, namely, $10^{6}\mathrm{\,M_{\odot}}$. We observe that the relative difference in binding energy exhibits converging behavior with respect to the $\bar{E}_{b}$ benchmark. Presuming that the largest radius provides the most accurate binding energy ($\bar{E}_{b}$), we find the relative difference between the binding energies at smallest outer boundary radius and the largest one is $\sim 0.3\%$. Furthermore, the relative difference in binding energies between $5\times 10^{5}\mathrm{\,M_{\odot}}$ and $1\times 10^{6}\mathrm{\,M_{\odot}}$ radii is $\sim 0.04\%$.
These results suggest that we can construct constraint satisfying and self-consistent ID for different values of the outer boundary radius of the grid.
5 Summary
In this work we have presented a significant upgrade to the Elliptica infrastructure. Previously limited to black hole-neutron star system and (piecewise) polytropic EOS for NS, Elliptica can now construct ID for spinning BNS systems, incorporating realistic tabulated EOS for the NS matter.
To incorporate tabular EOS, we have developed a number of techniques centered around interpolation that allow convergence in the resulting solution while remaining as true as possible to the underlying microphysics. While the convergence of the solution is negatively affected by the complexity of the EOS, we maintain convergence up to a limit imposed by the surface features.
In this work, we selected tabulated SLy4 and SFHo EOSs from the CompOSE tables. Since these EOS are generic among available cold NS EOS tables, our numerical methods remain still applicable to many such EOS. Additionally, our conclusions regarding convergence, for example, are relevant because the limiting factors on accuracy are common across these families of EOS.
We have validated our code through convergence tests and comparisons with established analytical results, particularly in the post-Newtonian regime. Additionally, by setting the outer boundary radius at different values, we study the convergence of the constraints and the binding energy of the system. We observe that the constraints exhibit spectral convergence in patches covering the large-radius regions of the grid. Furthermore, we find that choosing an outer boundary radius in the range of $(10^{5},10^{6})\mathrm{\,M_{\odot}}$ results in an uncertainty bound of less than $\sim 0.2\%$ in the binding energy of the BNS system. These tests demonstrate the code’s accuracy and robustness.
For smooth EOS, such as polytropic models, we achieve spectral convergence, indicating an optimal error reduction rate with increasing resolution. However, for EOS tables with discontinuities, spectral convergence is diminished. At high resolutions, we observe a convergence plateau inside NS, though global convergence is still ensured. This behavior is expected for discontinuous functions.
While we have not studied our ID in the context of an evolution simulation in this work, it is important to assess the quality of the ID after its interpolation onto the grid of an evolution code. In particular, the propagation of constraint violations and the oscillation of the rest mass density at the center of each NS are important diagnostics to be studied. Moreover, since different ID codes employ different numerical parameters and algorithms to solve ID equations, it is important to examine the effects of these differences on the resultant gravitational waves. Specifically, from the point of view of the next generation detectors such as Cosmic Explorer [11], we need to study the mismatch between two gravitational wave signals emitted from the evolution of a BNS system started from the same ID but using two different ID codes. These topics are left for future work.
While some first steps have been taken to construct self-consistent ID for magnetized NS by solving the Einstein-Euler-Maxwell equations assuming ideal magnetohydrodynamic [99, 100], there are no ID with self-consistent magnetic field for binary systems. As such, a particularly interesting area for future development lies in constructing self-consistent ID for BNS systems that contain magnetic fields. This capability would be specially valuable for studying systems containing pulsars or magnetars, where these fields play a significant role in the dynamics.
AR and AN gratefully acknowledge David Radice for valuable discussions and insightful feedback on the manuscript. AR acknowledges support from NASA under award No. 80NSSC21K1720. The numerical simulations were performed on Roar Collab High Performance Computing Cluster at The Pennsylvania State University. Elliptica is now public  [^3] https://github.com/rashti-alireza/Elliptica and open-source under GPL-3.0 license.
Appendix A Interpolation of the equation of state
To generate a suitable interpolant of a tabular EOS, we first approximate $\frac{\partial p}{\partial h}$, $\frac{\partial\epsilon}{\partial h}$, and $\frac{\partial\rho_{0}}{\partial h}$ by finite-difference methods on the CompOSE data using Fornberg’s algorithm [101], explained in B, for unevenly-spaced grids. Having the derivatives, we then generate a spline interpolant composed of Hermite polynomials. The procedure for e.g. the pressure $p\left(h\right)$ is:
Step 0: We are given data points $(p_{j},h_{j})$ monotonically increasing in $h$.
Step 1: For each point $j$, we take the set of $N$ points $X_{j}=\left\{h_{j-\frac{N}{2}},...,h_{j},...,h_{j+\frac{N}{2}}\right\}$ centered around $h_{j}$. We shift the indices when needed for points near the boundaries $h_{0}$ and $h_{j_{\mathrm{max}}}$.
Step 2: At each point, approximate $\left.\frac{dp}{dh}\right|_{h_{j}}\approx\sum_{\nu=j-\frac{N}{2}}^{j+\frac{N}{2}}\delta_{N,\nu}^{1}p_{\nu}$, with the coefficients $\delta$ calculated by Fornberg’s algorithm applied over the set $X_{j}$ obtained in Step 1.
Step 3: Having $p_{j}$ and its derivative $\frac{dp(h_{j})}{dh}$, we generate the interpolating Hermite polynomial (according to [102]) of desired order.
The same procedure is applied to obtain splines for $e\left(h\right)$, $\rho_{0}\left(h\right)$, and $\epsilon\left(h\right)$. Derivatives of these functions are thereafter approximated by analytical derivatives of the spline interpolant.
Appendix B Application of Fornberg’s finite difference method
Step 2 of Sec. (A) entails finding derivatives such as $\frac{dp}{dh}$, which EOS tables do not provide. We evaluate these derivatives numerically using finite difference methods on the data points $\left(p_{j},h_{j}\right)$. Since the data points are typically unevenly spaced, we use Fornberg’s finite difference algorithm to generate finite difference coefficients that approximate $\frac{dp(h_{j})}{dh}$ at each point [101]. Specifically, the algorithm calculates the weights $\delta_{N,\nu}^{m}$ such that
$\left.\frac{d^{m}p}{{dh}^{m}}\right|_{h_{j}}\approx\sum_{\nu=j-\frac{N}{2}}^{j+\frac{N}{2}}\delta_{N,\nu}^{m}p_{\nu},$ (57)
where $N$ is the number of points used (and determines the order of the finite difference approximation).
While Fornberg’s algorithm can provide derivatives of arbitrarily high orders (given enough data points), we are only interested in the first derivative. In addition, we do not need the finite difference coefficients for all orders, so we simplify the algorithm slightly. Then given a subset $X_{j}$ centered around $h_{j}$, from Step 1 of Sec. (A), we find the finite difference weights using algorithm 2 (where $c_{1}$, $c_{2}$, and $c_{3}$ are introduced just to simplify the notation, and $x_{k}$ is the $k$th element of $X_{j}$).
Algorithm 2 Simplification of Fornberg’s finite difference algorithm, adapted from [101].
$\delta^{0}_{0,0}=c_{1}=1$
2:for $n=1,...,N$ do
3:     $c_{2}=1$
4:     for $\nu=0,...,n-1$ do
5:         $c_{3}=x_{n}-x_{\nu}$
6:         $c_{2}=c_{2}\cdot c_{3}$
7:         if $n=0,1$ then
8:              $\delta^{n-1}_{n-1,\nu}=0$
9:         end if
10:         $\delta^{0}_{n,\nu}=\frac{\left(x_{n}-h_{j}\right)}{c_{3}}\delta^{0}_{n-1,\nu}$
11:         $\delta^{1}_{n,\nu}=\frac{1}{c_{3}}\left(\left(x_{n}-h_{j}\right)\delta^{1}_{n-1,\nu}-\delta^{0}_{n-1,\nu}\right)$
12:     end for
13:     $\delta^{0}_{n,n}=\frac{c_{1}}{c_{2}}\left(h_{j}-x_{n-1}\right)\delta^{0}_{n-1,n-1}$
14:     $\delta^{1}_{n,n}=\frac{c_{1}}{c_{2}}\left(\delta^{0}_{n-1,n-1}+\left(h_{j}-x_{n-1}\right)\delta^{1}_{n-1,n-1}\right)$
15:     $c_{1}=c_{2}$
16:end for